OpenStudy (anonymous):

How do I Find a unit vector in the direction of w=(-4,6)

OpenStudy (anonymous):

Do you want the magnitude and direction?

OpenStudy (anonymous):

Magnitude =\[\sqrt{y^2+x^2}\]

OpenStudy (anonymous):

direction

OpenStudy (turingtest):

to make you vector into a unit vector you divide it by the magnitude the direction will not change

OpenStudy (anonymous):

The angle would be between180 and 90 degrees (not sure what quadrant that is)

OpenStudy (turingtest):

your*

OpenStudy (anonymous):

so tan(4/6) equals direction, but I am not sure if tan is positive or negative in that quadrant.

OpenStudy (turingtest):

"How do I Find a unit vector in the direction of w=(-4,6)" doesn't ask for an angle, it just asks for the unit vector that points in the same direction, so just divide your original vector by the formula snipa gave you

OpenStudy (anonymous):

Well, a vector, by definition, has both magnitude AND direction.

OpenStudy (turingtest):

and it would be arctan(-2/3) for direction if you did want to know it

OpenStudy (anonymous):

Yeah, @TuringTest is right. It is arctan to find the angle. Sry

OpenStudy (turingtest):

yes I agree, so you don't need to specify the angle unless they ask for it

OpenStudy (turingtest):

no apologies necessary :) the point is that the unit vector is the original vector divided by the magnitude. Do you understand that @parksbap ?

OpenStudy (anonymous):

and tan(x) is positive in the first and second quadrants (between 0 and 180 degrees)

OpenStudy (anonymous):

yes and no

OpenStudy (anonymous):

a little confusing but im catching it

OpenStudy (anonymous):

The direction would be 146.3 degrees. I think

OpenStudy (turingtest):

do you know how to find the magnitude of this vector?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Loook at my first formula

OpenStudy (turingtest):

@Snipa420 please try not to just shoot out numerical answers; that's not the goal of this site.

OpenStudy (anonymous):

i agree, i need to be shown how to do this for a test

OpenStudy (turingtest):

but yes, @Snipa420 has given the correct formula for the magnitude

OpenStudy (anonymous):

ok first formula get it!

OpenStudy (anonymous):

@TuringTest She needs to see how to find the direction as well, I think. Is that right, @parksbap

OpenStudy (turingtest):

so you should be able to take that, plus what I told you, and get the answer first: what is the magnitude?

OpenStudy (anonymous):

yes so i did the magnitude, i got \[\sqrt{52}\]

OpenStudy (anonymous):

so since there is only a decimal i need to break that down

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

so from that i have 4\[\sqrt{13}\]

OpenStudy (turingtest):

\[\sqrt{52}=\sqrt{4\times13}=2\sqrt{13}\]

OpenStudy (anonymous):

It would be\[\sqrt{4}*\sqrt{13}\]

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

ok, so now you just divide the vector you have by that, and that is your unit vector we can discuss the direction in radians or degrees again after that if you want...

OpenStudy (anonymous):

well the answer is -2/\[\sqrt{13} \] and 3/\[\sqrt{13}\]

OpenStudy (turingtest):

right and if you still need direction that's given by\[\tan^{-1}(\frac yx)\] since the unit vector points in the same direction as the original, you can use the original x and y in the formula

OpenStudy (anonymous):

@TuringTest , arctan(y/x) does not give you the real direction. It gives you a number that you must interpret further.

OpenStudy (turingtest):

how so??

OpenStudy (anonymous):

Just take arctan(y/x). YOu get around 33 degrees.

OpenStudy (anonymous):

You must then observe that because the point is in the 2nd quadrant(-4, 6), that 33 degrees is not the right answer

OpenStudy (turingtest):

\[\tan\theta=\frac yx\implies\theta=\tan^{-1}(\frac yx)\]ah yeah you're right about the whole adding \(\pi n\) thing, must have slipped my mind my mistake

OpenStudy (anonymous):

No, you dont add Pi/n either.

OpenStudy (turingtest):

you add pi...

OpenStudy (anonymous):

How I got to my answer is I saw that it was in the second quadrant, and thus did 180 degrees - arctan(y/x), and got THE RIGHT ANSWER.

OpenStudy (anonymous):

Again, no you do not. If you add Pi, you will get an angle in the third quadrant.

OpenStudy (anonymous):

The angle we are looking for is in the second quadrant.

OpenStudy (turingtest):

I get -33 deg on my calc, so addin 180 puts me in QII

OpenStudy (anonymous):

Oh ok. YOu are right.

OpenStudy (turingtest):

\[\tan^{-1}(-2/3)\approx-33.69\]teamwork!

OpenStudy (anonymous):

I just took arctan(y/x), like you previously said, and then did 180-that answer.

OpenStudy (anonymous):

So we got the Same answer!!! Nice. My bad for doubting your maths skillz

OpenStudy (turingtest):

so yeah, just add pi why? because tan never changes when you add \(\pi n\) if n is an integer

OpenStudy (turingtest):

it's all good, my math skills may exist, but my fatigue always could use a second set of eyes thanks Snipa

OpenStudy (anonymous):

you are right, you are right...

OpenStudy (anonymous):

Yeah no problem. Later